# Weight Conditioning for Smooth Optimization of Neural Networks

### v1

@@ 1 Introduction
Normalization techniques, including batch normalization [8], weight standardization [14], and weight normalization [17], have become fundamental to the advancement of deep learning, playing a critical role in the development and performance optimization of deep learning models for vision applications [5, 7, 22, 11]. By ensuring consistent scales of inputs and internal activations, these methods not only stabilize and accelerate the convergence process but also mitigate issues such as vanishing or exploding gradients. In this paper, we put forth a normalization method termed weight conditioning, designed to help in the optimization of neural network architectures, including both feedforward and convolutional layers, through strategic manipulation of their weight matrices. By multiplying a weight matrix of a neural architecture by a predetermined matrix conditioner, weight conditioning aims to minimize the condition number of these weight matrices, effectively narrowing the disparity between their smallest and largest singular values. Our findings demonstrate that such conditioning not only influences the Hessian matrix of the associated loss function when training such networks, leading to a lower condition number but also significantly enhances the efficiency of iterative optimization methods like gradient descent by fostering a smoother loss landscape. Our theoretical analysis elucidates the impact of weight matrices on the Hessian’s condition number, revealing our central insight: optimizing the condition number of weight matrices directly facilitates Hessian conditioning, thereby expediting convergence in gradient-based optimization by allowing a larger learning rate to be used. Furthermore, weight conditioning can be used as a drop-in component along with the above normalization techniques, yielding better accuracy and training on a variety of problems.
@@ 2.0.1 Normalization in deep learning:
Normalization techniques have become pivotal in enhancing the training stability and performance of deep learning models. Batch normalization, introduced by Ioffe and Szegedy [8], normalizes the inputs across the batch to reduce internal covariate shift, significantly improving the training speed and stability of neural networks. Layer normalization, proposed by Ba et al. [1], extends this idea by normalizing inputs across features for each sample, proving particularly effective in recurrent neural network architectures [1] and transformer networks [21]. Weight normalization, by Salimans and Kingma [17], decouples the magnitude of the weights from their direction, facilitating a smoother optimization landscape. Lastly, weight standardization, introduced by Qiao et al. [14], standardizes the weights in convolutional layers, further aiding in the optimization process, especially when combined with batch normalization. Together, these techniques address various challenges in training deep learning models, underscoring the continuous evolution of strategies to improve model convergence and performance.
@@ 3 Notation
where the weights $W_{k}\in\mathbb{R}^{n_{k-1}\times n_{k}}$ and the biases $b_{k}\in\mathbb{R}^{n_{k}}$ and $\phi$ is an activation applied component wise. The notation $[m]$ is defined by $[m]=\{1,\ldots,m\}$. We will also fix a loss function $\mathcal{L}$ for minimizing the weights of $F$. In the experiments this will always be the MSE loss or the Binary Cross Entropy (BCE) loss. Note that $\mathcal{L}$ depends on $F$.
@@ 4 Motivation
This means we must have $\eta<\frac{2}{\sigma_{i}}$ for each $i$ and this will be satisfied if $\eta<\frac{2}{\sigma_{1}}$ since $\sigma_{1}$ is the largest singular value. Furthermore, we see that the progress of gradient descent in the $ith$ direction $v_{i}$ is bounded by
@@
Preconditioning involves the application of a matrix, known as the preconditioner $P$, to another matrix $A$, resulting in the product $PA$, with the aim of achieving $\kappa(PA)\leq\kappa(A)$ [13, 15]. This process, typically referred to as left preconditioning due to the multiplication of $A$ from the left by $P$, is an effective method to reduce the condition number of $A$. Besides left preconditioning, there is also right preconditioning, which considers the product $AP$, and double preconditioning, which employs two preconditioners, $P_{1}$ and $P_{2}$, to form $P_{1}AP_{2}$. Diagonal matrices are frequently chosen as preconditioners because their application involves scaling the rows or columns of $A$, thus minimally adding to the computational cost of the problem. Examples of preconditioners are:
@@
    Jacobi Preconditioner: Given a square matrix $A$ the Jacobi preconditioner $D$ consists of the inverse of the diagonal elements of $A$, $A\rightarrow diag(A)^{-1}A$ [9].
@@
    Row Equilibration: Given a $n\times m$ matrix $A$, row equilibration is a diagonal $n\times n$ matrix with the inverse of the 2-norm of each row of $A$ on the diagonal, $A\rightarrow(||A_{i:}||_{2})^{-1}A$, where $A_{i:}$ denotes the ith-row of $A$ [2].
@@
    Column Equilibration: Given a $n\times m$ matrix $A$, column equilibration is a diagonal $m\times m$ matrix with the inverse of the 2-norm of each column of $A$ on the diagonal, $A\rightarrow A(||A_{:i}||_{2})^{-1}$, where $A_{:i}$ denotes the ith-column of $A$.
@@
The interested reader can consult [13, 15] for more on preconditioner. In Sec. 5.1 we will explain how row equilibration helps reduce the condition number.
@@
~$\widetilde{\mathcal{L}}(\theta)\approx\frac{1}{2}(\theta-\theta_{0})^{T}H(\theta_{0})(\theta-\theta_{0})+(\nabla\widetilde{\mathcal{L}}(\theta_{0}))(\theta-\theta_{0})+\widetilde{\mathcal{L}}(\theta_{0})$ (14)
@@ 5.1 Weight conditioning for feed forward networks
Our objective is to rigorously demonstrate that an optimal choice for weight conditioning a network is to use row equilibration. Recall from sec. 4, that row equilibration preconditioner for a weight matrix $W_{k}\in\mathbb{R}^{n_{k-1}\times n_{k}}$ is defined as a diagonal matrix $E_{k}\in\mathbb{R}^{n_{k-1}\times n_{k-1}}$. Each diagonal element of $E_{k}$ is determined by the inverse of the 2-norm of the $i$-th row vector of $W_{k}$, i.e. $||(W_{k})_{i:}||_{2}^{-1}$.
@@
There are two main reasons we are choosing to use equilibration as the preferred form to condition the weights of the neural network $F$. The first reason is that it is a diagonal preconditioner, therefore requiring low computational cost to compute. Though more importantly, as the following theorem of Van Der Sluis [18] shows, it is the optimum preconditioner amongst all diagonal preconditioners to reduce the condition number.
Theorem 5.1 (Van Der Sluis [18]).
@@ 6.1 Convolutional Neural Networks (CNNs)
CNNs are pivotal for vision-related tasks, thriving in image classification challenges due to their specialized architecture that efficiently learns spatial feature hierarchies. Among the notable CNN architectures in the literature, we will specifically explore two significant ones for our experiment: the Inception [20] architecture and DenseNet [6], both of which continue to be highly relevant and influential in the field.
@@ 6.1.2 Inception:
In our experiment, we employed a modified Inception architecture, as proposed in [20], featuring eight inception blocks. Weight conditioning was effectively applied to just the initial convolutional layer and the final linear layer. For detailed insights into the architecture and normalization applications, see Sec. 2 of the Supp. material.
@@ 6.1.3 DenseNet:
For this experiment we tested four normalizations on the DenseNet architecture [6]. We found that it sufficed to apply equilibrated weight conditioning to the first convolutional layer and the last feedforward layer as in the case for Inception. An in-depth description of the architecture employed in our study, and how we applied each normalization is given in Sec. 2 of the Supp. material.
@@ 6.2 Vision Transformers (ViTs)
Vision Transformers (ViTs) [4] have emerged as innovative architectures in computer vision, showcasing exceptional capabilities across a broad spectrum of tasks. In general, we found that most work in the literature applied layer normalization to vision transformers which was much more robust for training than batch normalization. We found that if we removed layer normalization training was impeded significantly leading to gradients not being able to be backpropgated. Therefore, for this experiment we will consider 3 normalization scenarios, layer normalization (LN), layer normalization with weight normalization (LN + W) and layer normalization with equilibrated weight conditioning (LN + E).
@@ 6.2.1 ViT-Base (ViT-B):
The ViT-B architecture [19], with its 86 million parameters, exemplifies a highly overparameterized neural network. We investigated three variations of ViT-B, each modified with a different normalization: LN, LN + W, and LN + E; the implementation details are provided in Sec. 2 of the Supp. material. These models were trained on the ImageNet1k dataset using a batch size of 1024 and optimized with AdamW.
@@ 6.3 Neural Radiance Fields (NeRF)
Neural Radiance Fields (NeRF) [12, 3, 10, 16] have emerged as a pioneering approach in 3D modeling, using Multi-Layer Perceptrons (MLPs) to reconstruct 3D objects and scenes from multi-view 2D images. We utilized the standard NeRF model from [12], noting that unlike transformers or CNNs, NeRF architectures are relatively shallow, typically comprising 8 hidden layers, where common normalization techniques can hinder training. We explored the performance of a standard NeRF against an equilibrated weight conditioned NeRF (E-NeRF). For an in-depth look at the NeRF setup and our application of weight conditioning, refer to Sec. 2 in the Supp. material. Both models were trained on the LLFF dataset [12].
@@ Acknowledgements
Thomas X Wang acknowledges financial support provided by DL4CLIM (ANR- 19-CHIA-0018-01), DEEPNUM (ANR-21-CE23-0017-02), PHLUSIM (ANR-23-CE23- 0025-02), and PEPR Sharp (ANR-23-PEIA-0008”, “ANR”, “FRANCE 2030”).
Part of this work was performed using HPC resources from GENCI–IDRIS (Grant 2023-AD011013332R1).

### v2

@@ 1 Introduction
Normalization techniques, including batch normalization [ioffe2015batch], weight standardization [qiao2019micro], and weight normalization [salimans2016weight], have become fundamental to the advancement of deep learning, playing a critical role in the development and performance optimization of deep learning models for vision applications [he2016deep, huang2023normalization, zhang2020exemplar, lubana2021beyond]. By ensuring consistent scales of inputs and internal activations, these methods not only stabilize and accelerate the convergence process but also mitigate issues such as vanishing or exploding gradients. In this paper, we put forth a normalization method termed weight conditioning, designed to help in the optimization of neural network architectures, including both feedforward and convolutional layers, through strategic manipulation of their weight matrices. By multiplying a weight matrix of a neural architecture by a predetermined matrix conditioner, weight conditioning aims to minimize the condition number of these weight matrices, effectively narrowing the disparity between their smallest and largest singular values. Our findings demonstrate that such conditioning not only influences the Hessian matrix of the associated loss function when training such networks, leading to a lower condition number but also significantly enhances the efficiency of iterative optimization methods like gradient descent by fostering a smoother loss landscape. Our theoretical analysis elucidates the impact of weight matrices on the Hessian’s condition number, revealing our central insight: optimizing the condition number of weight matrices directly facilitates Hessian conditioning, thereby expediting convergence in gradient-based optimization by allowing a larger learning rate to be used. Furthermore, weight conditioning can be used as a drop-in component along with the above normalization techniques, yielding better accuracy and training on a variety of problems.
@@ 2.0.1 Normalization in deep learning:
Normalization techniques have become pivotal in enhancing the training stability and performance of deep learning models. Batch normalization, introduced by Ioffe and Szegedy [ioffe2015batch], normalizes the inputs across the batch to reduce internal covariate shift, significantly improving the training speed and stability of neural networks. Layer normalization, proposed by Ba et al. [ba2016layer], extends this idea by normalizing inputs across features for each sample, proving particularly effective in recurrent neural network architectures [ba2016layer] and transformer networks [xiong2020layer]. Weight normalization, by Salimans and Kingma [salimans2016weight], decouples the magnitude of the weights from their direction, facilitating a smoother optimization landscape. Lastly, weight standardization, introduced by Qiao et al. [qiao2019micro], standardizes the weights in convolutional layers, further aiding in the optimization process, especially when combined with batch normalization. Together, these techniques address various challenges in training deep learning models, underscoring the continuous evolution of strategies to improve model convergence and performance.
@@ 3 Notation
where the weights $W_{k}\in\mathbb{R}^{n_{k-1}\times n_{k}}$ and the biases $b_{k}\in\mathbb{R}^{n_{k}}$ and $\phi$ is an activation applied component wise. The notation $[m]$ is defined by $m]=\{1,\ldots,m\}$. We will also fix a loss function $\mathcal{L}$ for minimizing the weights of $F$. In the experiments this will always be the MSE loss or the Binary Cross Entropy (BCE) loss. Note that $\mathcal{L}$ depends on $F$.
@@ 4 Motivation
This means we must have $\eta<\frac{2}{\sigma_{i}}$ for each $i$ and this will be satisfied if $\eta<\frac{2}{\sigma_{1}}$ since $\sigma_{1}$ is the largest singular value. Furthermore, we see that the progress of gradient descent in the ithith direction $v_{i}$ is bounded by
@@
Preconditioning involves the application of a matrix, known as the preconditioner $P$, to another matrix $A$, resulting in the product $PA$, with the aim of achieving $\kappa(PA)\leq\kappa(A)$ [nocedal1999numerical, qu2022optimal]. This process, typically referred to as left preconditioning due to the multiplication of $A$ from the left by $P$, is an effective method to reduce the condition number of $A$. Besides left preconditioning, there is also right preconditioning, which considers the product $AP$, and double preconditioning, which employs two preconditioners, $P_{1}$ and $P_{2}$, to form $P_{1}AP_{2}$. Diagonal matrices are frequently chosen as preconditioners because their application involves scaling the rows or columns of $A$, thus minimally adding to the computational cost of the problem. Examples of preconditioners are:
@@
    Jacobi Preconditioner: Given a square matrix $A$ the Jacobi preconditioner $D$ consists of the inverse of the diagonal elements of $A$, $A\rightarrow diag(A)^{-1}A$ [jacobi1845ueber].
@@
[…]
@@
[…]
@@
The interested reader can consult [nocedal1999numerical, qu2022optimal] for more on preconditioner. In Sec. 5.1 we will explain how row equilibration helps reduce the condition number.
@@
$\widetilde{\mathcal{L}}(\theta)\approx\frac{1}{2}(\theta-\theta_{0})^{T}H(\theta_{0})(\theta-\theta_{0})+(\nabla\widetilde{\mathcal{L}}(\theta_{0}))(\theta-\theta_{0})+\widetilde{\mathcal{L}}(\theta_{0})$ (14)
@@ 5.1 Weight conditioning for feed forward networks
Our objective is to rigorously demonstrate that an optimal choice for weight conditioning a network is to use row equilibration. Recall from sec. 4, that row equilibration preconditioner for a weight matrix $W_{k}\in\mathbb{R}^{n_{k-1}\times n_{k}}$ is defined as a diagonal matrix $E_{k}\in\mathbb{R}^{n_{k-1}\times n_{k-1}}$. Each diagonal element of $E_{k}$ is determined by the inverse of the 2-norm of the $i$-th row vector of $W_{k}$, i.e. $(W_{k})_{i:}||_{2}^{-1}$.
@@
There are two main reasons we are choosing to use equilibration as the preferred form to condition the weights of the neural network $F$. The first reason is that it is a diagonal preconditioner, therefore requiring low computational cost to compute. Though more importantly, as the following theorem of Van Der Sluis [van1969condition] shows, it is the optimum preconditioner amongst all diagonal preconditioners to reduce the condition number.
Theorem 5.1(Van Der Sluis [van1969condition]).
@@ 6.1 Convolutional Neural Networks (CNNs)
CNNs are pivotal for vision-related tasks, thriving in image classification challenges due to their specialized architecture that efficiently learns spatial feature hierarchies. Among the notable CNN architectures in the literature, we will specifically explore two significant ones for our experiment: the Inception [szegedy2015going] architecture and DenseNet [huang2017densely], both of which continue to be highly relevant and influential in the field.
@@ 6.1.2 Inception:
In our experiment, we employed a modified Inception architecture, as proposed in [szegedy2015going], featuring eight inception blocks. Weight conditioning was effectively applied to just the initial convolutional layer and the final linear layer. For detailed insights into the architecture and normalization applications, see Sec. 2 of the Supp. material.
@@ 6.1.3 DenseNet:
For this experiment we tested four normalizations on the DenseNet architecture [huang2017densely]. We found that it sufficed to apply equilibrated weight conditioning to the first convolutional layer and the last feedforward layer as in the case for Inception. An in-depth description of the architecture employed in our study, and how we applied each normalization is given in Sec. 2 of the Supp. material.
@@ 6.2 Vision Transformers (ViTs)
Vision Transformers (ViTs) [dosovitskiy2020image] have emerged as innovative architectures in computer vision, showcasing exceptional capabilities across a broad spectrum of tasks. In general, we found that most work in the literature applied layer normalization to vision transformers which was much more robust for training than batch normalization. We found that if we removed layer normalization training was impeded significantly leading to gradients not being able to be backpropgated. Therefore, for this experiment we will consider 3 normalization scenarios, layer normalization (LN), layer normalization with weight normalization (LN + W) and layer normalization with equilibrated weight conditioning (LN + E).
@@ 6.2.1 ViT-Base (ViT-B):
The ViT-B architecture [steiner2021train], with its 86 million parameters, exemplifies a highly overparameterized neural network. We investigated three variations of ViT-B, each modified with a different normalization: LN, LN + W, and LN + E; the implementation details are provided in Sec. 2 of the Supp. material. These models were trained on the ImageNet1k dataset using a batch size of 1024 and optimized with AdamW.
@@ 6.3 Neural Radiance Fields (NeRF)
Neural Radiance Fields (NeRF) [nerf, chng2022gaussian, lin2021barf, reiser2021kilonerf] have emerged as a pioneering approach in 3D modeling, using Multi-Layer Perceptrons (MLPs) to reconstruct 3D objects and scenes from multi-view 2D images. We utilized the standard NeRF model from [nerf], noting that unlike transformers or CNNs, NeRF architectures are relatively shallow, typically comprising 8 hidden layers, where common normalization techniques can hinder training. We explored the performance of a standard NeRF against an equilibrated weight conditioned NeRF (E-NeRF). For an in-depth look at the NeRF setup and our application of weight conditioning, refer to Sec. 2 in the Supp. material. Both models were trained on the LLFF dataset [nerf].
@@ Acknowledgements
Simon Lucey acknowledges support from the Australian Research Council (ARC) through the Discovery Project DP220103803. Thomas X Wang acknowledges financial support provided by DL4CLIM (ANR- 19-CHIA-0018-01), DEEPNUM (ANR-21-CE23-0017-02), PHLUSIM (ANR-23-CE23- 0025-02), and PEPR Sharp (ANR-23-PEIA-0008”, “ANR”, “FRANCE 2030”). Part of this work was performed using HPC resources from GENCI–IDRIS (Grant 2023-AD011013332R1).